OpenStudy (anonymous):

use the lim f(x)-f(a)/(x-a) to find the derivative of 5x^2 - 3x + 2, a =1 x->a

OpenStudy (anonymous):

the derivative equation is \[\lim_{h \rightarrow 0}\frac{f(x+h)-f(x)}{h}\] try using that instead of wahtever you have there

OpenStudy (anonymous):

I realize this, but on my quiz we have to know how to use the other one, because we are looking for a specific point. I understand you substitute, but I am having problems with the factoring. I am getting (5x^2 -3x + 23)/(x-1)

OpenStudy (anonymous):

It's called the definition of a derivative at a point.

OpenStudy (anonymous):

oh i see what an odd equation, but it makes perfect sense unless you can factor 5x^2...., you're done

OpenStudy (anonymous):

Oh. Thought there was more factoring to be done. Thanks.

OpenStudy (anonymous):

oh wait hold that thought

OpenStudy (anonymous):

ok big issue, you need to find the limit as x->1 this will give you a zero on the bottom so what you do is you divide all the terms by x and simplify it then you find the limit

OpenStudy (anonymous):

The simplifying is where I'm having the problem. I'm stuck with that fraction and don't know where to go from there.

OpenStudy (anonymous):

\[\lim_{x \rightarrow a}\frac{5x^2 -3x + 23}{x-1}\]

OpenStudy (anonymous):

actually, no, im completely clueless a =1 not infinity or 0, theres not way to get rid of the x-1 in order to find the limit

OpenStudy (anonymous):

Exactly. Any ideas? Did I do something wrong to get that fraction?

OpenStudy (anonymous):

how did you get 23? but that doesnt really change anything

OpenStudy (anonymous):

plug a = 1 into the function, you get 24. When you write the fraction after substitution like thid \[\lim_{x \rightarrow 1}\frac{ 5x ^{2}-3x+24-1 }{ x-1 }\]

OpenStudy (anonymous):

24-1=23

OpenStudy (anonymous):

the formula says f(x)-f(a), so that is why it is -1

OpenStudy (anonymous):

\[\lim_{x \rightarrow a} \frac{5x^2 - 3x + 2-(5a^2 - 3a + 2)}{x-a}\] \[\lim_{x \rightarrow a} \frac{5(x^2-a^2) - 3(x-a)}{x-a}\] \[\lim_{x \rightarrow a} \frac{5(x-a)(x+a) - 3(x-a)}{x-a}\] \[\lim_{x \rightarrow a} \frac{(x-a)(5(x+a) - 3)}{x-a}\] \[\lim_{x \rightarrow a} 5(x+a) - 3\] \[5(a+a)-3\] a=1 7

OpenStudy (anonymous):

substitution always comes last my mistake

OpenStudy (anonymous):

and this is why i cant be lazy and rely off other people's work :(

OpenStudy (anonymous):

This looks likes the same process as the other formula. Is it essentially?

OpenStudy (anonymous):

it is the same thing, sort of basically you need to understand the slope equation

OpenStudy (anonymous):

[f(x+h)-f(x)]/(x+h-x) in this case, x+h=x and x is equal to a

OpenStudy (anonymous):

Okay, thanks for your help.